OpenStudy (anonymous):

What is the y coordinate of the y−intercept of the line that passes through the points (−3, −1) and (1, 7)? I really need help this is the last question on my assignment!

OpenStudy (shamim):

formula \[y-y _{1}=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }(x-x _{1})\]is

OpenStudy (shamim):

here\[x _{1}=-3,y _{1}=-1,x _{2}=1,y _{2}=7\]

OpenStudy (shamim):

the slope of ur line is\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (shamim):

now try to find out the equation of line

OpenStudy (anonymous):

Oh i get it! thanks so much !(:

OpenStudy (shamim):

welcome

OpenStudy (shamim):

but u hv to find out y intercept

OpenStudy (shamim):

did u find out y intercept

OpenStudy (shamim):

anyway wts the equation of line u got

OpenStudy (anonymous):

First we need to find the slope .... slope(m) = (y2 - y1)/(x2 - x1) set 1 (-3,-1) set 2 (1,7) now sub in your points m = (7 - (-1)/ (1 - (-3) m = (7 + 1)/(1 + 3) m = 8/4 = 2 your slope is 2 now we will use this formula : y = mx + b using slope 2 and either of your points....I will use (1,7) y = mx + b 7 = 2(1) + b 7 = 2 + b 7 - 2 = b 5 = b your equation in slope intercept form is y = 2x + 5 If you need it in standard form,Ax + By = C, it is : y = 2x + 5 -2x + y = 5 (multiply by -1 to make x positive) 2x + (- y )= - 5